# Noisy Quantum Learning Theory

### v1

@@ Acknowledgments
The authors thank Sitan Chen, Soonwon Choi, Harald Putterman, and Ruohan Shen for valuable discussions. JC is supported by an Alfred P. Sloan Foundation Fellowship. WG is supported by NSF Grant CCF-2430375.
@@
In Appendix A, we review salient prior work on quantum sensing, learning, and computation in noisy settings. In Appendix B, we recall the definitions of several hybrid, relativized models of quantum computation used in our complexity-theoretic separations, and define the complexity class NBQP. In Appendix C.1, we review technical preliminaries on quantum information theory as well as quantum learning lower bounds. In Appendix D, we demonstrate our two superpolynomial oracle separations, proving Theorems 2.2 and 2.3, which comprise the complexity-theoretic portion of our results. In Appendix E, we demonstrate that the well-studied exponential quantum speedup for purity testing completely degrades under local noise, proving Theorem 2.4. We then turn to a physically-motivated reformulation of the purity-testing problem, demonstrating that the breakdown we showed previously can be rectified for the task of detecting a black hole microstate in the bulk of a tensor-network model for holographic duality, since the system has intrinsic error-correction properties. In Appendix F, we address the degradation of Pauli shadow tomography with noisy experiments, proving Theorem 2.6. In particular, we prove tight sample complexity lower bounds under three models of noisy quantum experiments, and provide noisy single-copy and two-copy algorithms for the task. In Appendix G, we prove several technical lemmas stated in earlier sections.
@@ Appendix A Related Work
Quantum sensing enhanced by quantum information processing
@@
Quantum sensing has a natural interpretation in the language of noisy quantum learning theory, since sensing targets are often Hamiltonian coefficients (e.g. field amplitudes or phases) or observables of quantum states (such as order parameters of quantum materials). In a future fault-tolerant setting, the most general sensing protocols could use a sensor to couple the experimental system, viewed as an oracle, to a quantum computer, thereby providing oracle access for an NBQP computation. Many works have established Heisenberg-limited (HL) scaling for traditional sensing methods such as Ramsey spectroscopy and interferometry in idealized settings [DLP03, DMMRD03, GLM04, KBD04]. However, under standard noise models, such as Markovian reservoirs or photon loss channels, the HL asymptotic scaling of these methods often reverts to the standard quantum limit (SQL) [YE04, BLFC07, MP07, SKHDD16, HSK+18, JWBA23].
@@
Quantum information processing has been used to either recover HL sensitivity in noisy environments or surpass HL in ideal settings. Quantum error correction on sensor qubits, introduced in [KLSL14], underlies much of the first direction. Subsequent work has developed error-correction protocols that recover asymptotic HL scaling for particular noise channels, such as dephasing or bosonic loss, and in specific architectures such as trapped-ion sensors [RSZM17, LZCJ19]. Ref. [ZZPJ18] provides a necessary and sufficient condition for such recoverability, showing that when the noise operators span the observables of interest, HL scaling typically cannot be restored and the protocol reverts to the SQL. In the NBQP model, we therefore expect many error-corrected sensing protocols to fail to retain HL asymptotics, since e.g. the depolarizing channel has a Kraus decomposition consisting of all $4^{n}$ Pauli strings. Practically, error-mitigation and approximate error-correction techniques remain promising for metrological gains despite asymptotic limitations [JCH14, ZJ20, RATG20], and careful co-design of error-correcting codes and sensing architectures with well-characterized noise can still enable HL sensing.
@@
Sensing beyond the HL does not yet fit into a single unified framework. A valuable non-entanglement resource in some beyond-HL proposals is quadrature squeezing: by compressing uncertainty in a relevant quadrature while enlarging it in an irrelevant one, and measuring only the former, interferometric protocols can achieve super-HL scaling in idealized regimes [PS08, GDD+13, OCW+24]. More recently, [AMC+25] combined Grover-type quantum speedups with sensing by searching for an ambient signal over discrete frequency bins, giving another beyond-HL framework that leverages deep quantum circuits. This work engineers a Grover phase oracle by wrapping the unknown signal with a quantum signal processing transform [MRTC21]. In the NBQP noise model where calls to this oracle are interleaved with noise, this speedup breaks down due to known results regarding Grover’s algorithm with noisy oracles [ABNR13, Ros23, Ros24]. Refining this algorithm for practical applications will require novel error-mitigation strategies both within the oracle construction and surrounding the oracle queries. An important practical follow-up to our work is to connect our noise-aware analysis of learning from uncharacterized systems with these beyond-HL protocols, which are largely developed only in ideal settings.
Noise-robust quantum learning
@@
In a related vein, the classical shadows algorithm [HKP20] with the standard Clifford ensemble is not noise-robust: [KG22] shows that under product depolarizing and amplitude-damping noise, its sample complexity for estimating Pauli operators scales exponentially with operator weight. By contrast, randomized measurement protocols using only single-qubit local control (and hence generating no entanglement) are expected to be significantly more noise-robust, since corruption of a constant fraction of qubits remains localized. Examples include quantum overlapping tomography [CW20] and classical shadows with a unitary 1-design ensemble.
@@
Hamiltonian learning from time dynamics is another quantum learning task where noise robustness is essential and closely tied to quantum metrology. Several works have proposed algorithms achieving $O(1/\varepsilon)$, Heisenberg-limited scaling in time complexity that are robust to state-preparation and measurement (SPAM) errors [HTFS23, HMG+25]. However, these protocols require deep circuits that interleave many layers of quantum control with queries to the unknown Hamiltonian. Moreover, [HMG+25] proves that such deep quantum control is necessary for Heisenberg-limited, ansatz-free Hamiltonian learning. Under interstitial noise, accumulated errors then destroy coherence unless the Hamiltonian coincidentally acts on the logical codespace of an error-correcting code. In a different direction, [CCH25] introduces quantum probe tomography, where a constant number of probes couple to a few sites of a large system to extract its parent Hamiltonian. While restricted to structured Hamiltonian classes, this probe setting requires neither probe entanglement nor deep control, yielding an end-to-end noise-robust, Hamiltonian learning strategy. However, it is not yet known if there are versions of quantum probe tomography which achieve the Heisenberg limit. In any case, such protocols highlight that it is prudent for practical Hamiltonian learning protocols to leverage the substantial structural constraints of physical Hamiltonians.
@@
Quantum learning in bosonic continuous-variable systems has also recently attracted attention. Bosonic statistics allow squeezing operations that reduce canonical quadrature variance below the standard Heisenberg limit. Leveraging this, [FIL+25] provides an algorithm for learning bosonic Gaussian unitaries, while [OCW+24] studies Gaussian channels; both algorithms exhibit a sample complexity that shrinks with the amount of squeezing used to prepare input states. Moreover, [OCW+24] demonstrates that the effects of realistic photon-loss, measurement, and crosstalk errors on estimation uncertainty are suppressed by a factor that grows exponentially with the squeezing parameter. Beyond such examples, the broader role of non-entanglement resources in quantum learning remains poorly understood.
@@
Even without formal noise-robustness guarantees, small-scale experiments have demonstrated quantum learning-enhanced efficiency. For instance, [CCL+19] use a strategy based on SWAP tests to estimate low-temperature properties of a Bose-Hubbard model in an optical lattice. To measure local observables at low temperatures, one needs only perform a variant of the SWAP test on a constant number of sites. However, if the target low temperature is separated from the physical temperature by a phase transition, or if one wishes to access the phase transition directly, then one may need to perform a version of the SWAP test on a number of sites that scales with $n$. Our exponential lower bound on noisy two-copy purity testing in Appendix E implicitly shows that such general SWAP-test-based strategies are exponentially hampered by local errors. We therefore expect a similar degradation for quantum virtual cooling when virtually cooling to at or below a phase transition, although the details will depend on the particular observables being measured and on whether the state-observable pair is intrinsically noise robust. To elaborate on this last point, it is possible that certain kinds of (macroscopic) observables are insensitive to certain kinds of local perturbations or errors, allowing reliable learning despite noise at every logical layer of a given quantum learning protocol. Clarifying the relationship between experimental learnability and the robustness of (macroscopic) observables to local errors remains an open question.
@@
Finally, [HFP22] develops a theoretical framework for quantum learning with noisy quantum devices, in a setting distinct from ours. They study noisy device learning from noisy access, whereas we consider learning uncharacterized and noisy systems with a fault-tolerant quantum computer. Their work analyzes a task similar to the Dec-IP problem from Definition F.2, which we use to establish quantum advantage with noisy access. However, [HFP22] considers access to perfect copies of the unknown state and diamond-error noise in the action of each gate for the upper bound; moreover, the given lower bounds neglect noise.
Quantum computation with noisy oracles
@@
The NBQP model also characterizes quantum computational problems making calls to noisy oracles, beyond the context of learning from quantum experiments. Ref. [CSS15] shows that a modification of the Bernstein-Vazirani problem retains a quantum speedup when only the output of the oracle is corrupted by depolarizing noise; it is simple to see that, under interstitial noise, this speedup vanishes.
@@
Several works study Grover’s search algorithm [Gro96] with various models of oracle noise, including phase error [SBW03], simple oracle call failure [RS08, ABNR13], and phase inversions [LLZT00], all finding that the time complexity under constant-strength noise reverts to asymptotically linear in the database size. Furthermore, [CCHL23] demonstrates such a slowdown for NISQ algorithms. Most aligned with our work, [Ros23] shows that if Grover oracle calls are sandwiched between layers of constant-strength global depolarizing or dephasing noise with all other operations perfect (a stronger computational model than NBQP), then the speedup once again breaks down, and in [Ros24] the result is strengthened to hold when depolarizing noise is applied only to a single qubit before and after the oracle query. While Grover oracles for database search tasks may eventually be implemented fault-tolerantly, these results suggest that achieving asymptotic metrological gains from the recently proposed Grover-enhanced sensing strategy in [AMC+25] may be difficult when ambient noise acts on the sensor qubits, effectively implementing an error channel before each oracle query.
@@ Appendix B Definitions
For the definition of the NISQ complexity class, we will need to make reference to a noise model for quantum computation. The most convenient is constant depolarizing noise per qubit, although as emphasized in [CCHL23] other noise models are suitable as well. To fix notation, a single-qubit depolarizing channel will be defined as follows.
@@
In the NISQ model, each use of the quantum device consists of a single noisy circuit call $\textnormal{NQC}^{O}_{\lambda}(n^{\prime},\{U\})$ followed by a complete measurement of all qubits in the computational basis; the only memory across calls is classical. In particular, the algorithm cannot perform intermediate quantum measurements, reset a subset of qubits, or introduce fresh clean ancillas in the middle of a coherent computation. Quantum error-correction and fault-tolerance schemes rely precisely on such operations: one repeatedly measures stabilizers, discards or resets noisy ancillas, and thereby pumps entropy out of the encoded state. Since none of this is available in the NISQ model, one can prove (see e.g. [CCHL23]) that the effective noise on the quantum state cannot be suppressed over time, and the class is inherently incapable of implementing full fault-tolerant quantum error correction.
@@ C.1 Quantum information theory toolkit
Lemma C.2 (Simulating arbitrary POVMs with rank-1 POVMs, e.g. Lemma 4.8 in [CCHL22]).
@@
Lemma C.5 ($\mathcal{M}_{k}^{cn}$ POVMs are $c$-query $k$-qubit circuits, Section 8.1 in [CGY24]).
@@ C.2 Tree representations for learning lower bounds
A powerful tool for proving information-theoretic lower bounds on the sample complexity of quantum learning tasks is the tree representation of a quantum learning algorithm. In this work, we bring together the learning tree formalisms proposed in previous works for memoryless algorithms, bounded-memory algorithms with limited memory lifetime, and bounded-memory algorithms with unbounded coherent lifetime [CGY24, CCHL22, CCHL23]. These frameworks fall under the following definition.
@@
We will leverage learning trees to prove query lower bounds for property testing problems (such as the many-vs. one distinguishing problem) using Le Cam’s two point method [Yu97]. The potential outcomes of a quantum learning algorithm for many-vs.-one distinguishing will be stored in classical memory, and encoded in the leaf nodes of the learning tree $\mathcal{T}$. Some leaves will correspond to outputting the maximally-mixed hypothesis, while others will guess the alternative event. If the two distributions over leaves $\ell$, namely $p_{\rho_{x}}(\ell)$ (given $\rho_{x}$ sampled from $\mathcal{D}$) and $p_{\mathds{1}/2^{n}}(\ell)$ (given the maximally mixed state) are very close to one another, our learning algorithm cannot successfully distinguish the two hypotheses with high accuracy. Formally:
@@
For a review of information-theoretic lower bound techniques including Le Cam’s method, see e.g. [Yu97] and its application to learning trees in [CCHL22]. Our goal will be to argue that for the total variation bound in Lemma C.11 to become appreciably large, the tree depth $T$ must grow exponentially in $n$. To do so, it suffices to bound the one-sided likelihood ratio. Let us first define this ratio, and then subsequently explain why bounding it is useful for us.
@@
    (Lemma 5.4 in [CCHL22]) If $L(\ell)\geq 1-\delta$ for all $\ell\in\textnormal{leaf}(\mathcal{T})$, then we have the bound $d_{\text{\rm TV}}(\mathbb{E}_{\mathcal{D}}[p_{\rho_{x}}(\ell)],p_{\mathds{1}/2^{n}}(\ell))\leq\delta$.
@@
    (Lemma 7 in [CGY24]) If the condition
@@
    Lemma 6 in [CGY24] For any $\delta\in(0,1)$,
@@
The second lemma tells us that if the likelihood ratio induced by intermediate nodes of the tree sharply concentrates (indicating that a single experiment carries little information about the unknown state), the ratio over leaves is likely close to $1$, and the tree depth must then be large to distinguish the hypotheses. Finally, our third lemma allows us to bound the total variation distance by a probabilistic leaf likelihood ratio. That is, if we consider a fixed but arbitrary root-to-leaf path and argue that, with high probability the likelihood ratio on that path is close to $1$, we can utilize Lemma C.11. This approach is useful when the bound on intermediate nodes fails, which may happen when a small, nonzero number of intermediate nodes has large likelihood ratio fluctuations despite only comprising a vanishing fraction of the eventual outcomes. For a detailed overview of these “edge-based” and “path-based” approaches, see [CCHL22].

### v2

@@ Acknowledgments
The authors thank Sitan Chen, Soonwon Choi, Harald Putterman, Ruohan Shen and Nikita Romanov for valuable discussions. JC is supported by an Alfred P. Sloan Foundation Fellowship. WG is supported by NSF Grant CCF-2430375. The authors do not declare any competing interests.
@@
In Appendix A, we review salient prior work on quantum sensing, learning, and computation in noisy settings. In Appendix B, we recall the definitions of several hybrid, relativized models of quantum computation used in our complexity-theoretic separations, and define the complexity class NBQP. In Appendix C.1, we review technical preliminaries on quantum information theory as well as quantum learning lower bounds. In Appendix D, we demonstrate our two superpolynomial oracle separations, proving Theorems 2.2 and 2.3, which comprise the complexity-theoretic portion of our results. In Appendix E, we demonstrate that the well-studied exponential quantum speedup for purity testing completely degrades under local noise, proving Theorem 2.4. We then turn to a physically-motivated reformulation of the purity-testing problem, demonstrating that the breakdown we showed previously can be rectified for the task of detecting a black hole microstate in the bulk of a tensor-network model for holographic duality, since the system has intrinsic error-correction properties. In Appendix F, we address the degradation of Pauli shadow tomography with noisy experiments, proving Theorem 2.6. In particular, we prove tight sample complexity lower bounds under three models of noisy quantum experiments, and provide noisy single-copy and two-copy algorithms for the task. In Appendix H, we prove several technical lemmas stated in earlier sections.
@@ Appendix A Related Work
Quantum sensing enhanced by quantum information processing.
@@
Quantum sensing has a natural interpretation in the language of noisy quantum learning theory, since sensing targets are often Hamiltonian coefficients (e.g. field amplitudes or phases) or observables of quantum states (such as order parameters of quantum materials). In a future fault-tolerant setting, the most general sensing protocols could use a sensor to couple the experimental system, viewed as an oracle, to a quantum computer, thereby providing oracle access for an NBQP computation. Many works have established Heisenberg-limited (HL) scaling for traditional sensing methods such as Ramsey spectroscopy and interferometry in idealized settings [D_Ariano_2003, De_Martini_2003, Giovannetti_2004, Kok_2004]. However, under standard noise models, such as Markovian reservoirs or photon loss channels, the HL asymptotic scaling of these methods often reverts to the standard quantum limit (SQL) [Yu_2004, Bellomo_2007, Monras_2007, Smirne_2016, Haase_2018, Jiao_2023].
@@
Quantum information processing has been used to either recover HL sensitivity in noisy environments or surpass HL in ideal settings. Quantum error correction on sensor qubits, introduced in [Kessler_2014], underlies much of the first direction. Subsequent work has developed error-correction protocols that recover asymptotic HL scaling for particular noise channels, such as dephasing or bosonic loss, and in specific architectures such as trapped-ion sensors [2017NatCo...8.1822R, Layden_2019]. Ref. [Zhou_2018] provides a necessary and sufficient condition for such recoverability, showing that when the noise operators span the observables of interest, HL scaling typically cannot be restored and the protocol reverts to the SQL. In the NBQP model, we therefore expect many error-corrected sensing protocols to fail to retain HL asymptotics, since e.g. the depolarizing channel has a Kraus decomposition consisting of all $4^{n}$ Pauli strings. This intuition is captured in our Theorem G.2, in which the well-known error-correction-based metrology scheme from [Zhou_2018] is degraded to the SQL beyond a sensing time threshold. A concurrent work [sahu2026achievingheisenberglimitusing] demonstrates that for sufficiently structured signals and noise channels, quantum error correction can recover HL sensitivity for noise rates below a threshold. In such specialized settings, our noisy HL sensing window can be extended indefinitely, at the expense of introducing a new threshold condition on the error rate. Practically, error-mitigation and approximate error-correction techniques remain promising for metrological gains despite asymptotic limitations [Jeske_2014, Zhou_2020, Rossi_2020], and careful co-design of error-correcting codes and sensing architectures with well-characterized noise can still enable HL sensing.
@@
Sensing beyond the HL does not yet fit into a single unified framework. A valuable non-entanglement resource in some beyond-HL proposals is quadrature squeezing: by compressing uncertainty in a relevant quadrature while enlarging it in an irrelevant one, and measuring only the former, interferometric protocols can achieve super-HL scaling in idealized regimes [Pezz__2008, Grote_2013, Oh_2024]. More recently, [allen2025quantumcomputingenhancedsensing] combined Grover-type quantum speedups with sensing by searching for an ambient signal over discrete frequency bins, giving another beyond-HL framework that leverages deep quantum circuits. This work engineers a Grover phase oracle by wrapping the unknown signal with a quantum signal processing transform [Martyn_2021]. In the NBQP noise model where calls to this oracle are interleaved with noise, this speedup breaks down due to known results regarding Grover’s algorithm with noisy oracles [ambainis_grover, rosmanis2023quantumsearchnoisyoracle, rosmanis2024addendumquantumsearchnoisy]. Refining this algorithm for practical applications will require novel error-mitigation strategies both within the oracle construction and surrounding the oracle queries. An important practical follow-up to our work is to connect our noise-aware analysis of learning from uncharacterized systems with these beyond-HL protocols, which are largely developed only in ideal settings.
Noise-robust quantum learning.
@@
In a related vein, the classical shadows algorithm [classical_shadow] with the standard Clifford ensemble is not noise-robust: [Koh_2022] shows that under product depolarizing and amplitude-damping noise, its sample complexity for estimating Pauli operators scales exponentially with operator weight. By contrast, randomized measurement protocols using only single-qubit local control (and hence generating no entanglement) are expected to be significantly more noise-robust, since corruption of a constant fraction of qubits remains localized. Examples include quantum overlapping tomography [cotler2020quantum] and classical shadows with a unitary 1-design ensemble.
@@
Hamiltonian learning from time dynamics is another quantum learning task where noise robustness is essential and closely tied to quantum metrology. Several works have proposed algorithms achieving $O(1/\varepsilon)$, Heisenberg-limited scaling in time complexity that are robust to state-preparation and measurement (SPAM) errors [Huang_2023, Hu_2025]. However, these protocols require deep circuits that interleave many layers of quantum control with queries to the unknown Hamiltonian. Moreover, [Hu_2025] proves that such deep quantum control is necessary for Heisenberg-limited, ansatz-free Hamiltonian learning. Under interstitial noise, accumulated errors then destroy coherence unless the Hamiltonian coincidentally acts on the logical codespace of an error-correcting code. In a different direction, [chen2025quantumprobetomography] introduces quantum probe tomography, where a constant number of probes couple to a few sites of a large system to extract its parent Hamiltonian. While restricted to structured Hamiltonian classes, this probe setting requires neither probe entanglement nor deep control, yielding an end-to-end noise-robust, Hamiltonian learning strategy. However, it is not yet known if there are versions of quantum probe tomography which achieve the Heisenberg limit. In any case, such protocols highlight that it is prudent for practical Hamiltonian learning protocols to leverage the substantial structural constraints of physical Hamiltonians.
@@
Quantum learning in bosonic continuous-variable systems has also recently attracted attention. Bosonic statistics allow squeezing operations that reduce canonical quadrature variance below the standard Heisenberg limit. Leveraging this, [fanizza2025efficientlearningbosonicgaussian] provides an algorithm for learning bosonic Gaussian unitaries, while [Oh_2024] studies bosonic random displacement channels; both algorithms exhibit a sample complexity that shrinks with the amount of squeezing used to prepare input states. Moreover, [Oh_2024] demonstrates that the effects of realistic photon-loss, measurement, and crosstalk errors on estimation uncertainty are suppressed by a factor that grows exponentially with the squeezing parameter. Beyond such examples, the broader role of non-entanglement resources in quantum learning remains poorly understood.
@@
Even without formal noise-robustness guarantees, small-scale experiments have demonstrated quantum learning-enhanced efficiency. For instance, [Cotler_2019] use a strategy based on SWAP tests to estimate low-temperature properties of a Bose-Hubbard model in an optical lattice. To measure local observables at low temperatures, one needs only perform a variant of the SWAP test on a constant number of sites. However, if the target low temperature is separated from the physical temperature by a phase transition, or if one wishes to access the phase transition directly, then one may need to perform a version of the SWAP test on a number of sites that scales with $n$. Our exponential lower bound on noisy two-copy purity testing in Appendix E implicitly shows that such general SWAP-test-based strategies are exponentially hampered by local errors. We therefore expect a similar degradation for quantum virtual cooling when virtually cooling to at or below a phase transition, although the details will depend on the particular observables being measured and on whether the state-observable pair is intrinsically noise robust. To elaborate on this last point, it is possible that certain kinds of (macroscopic) observables are insensitive to certain kinds of local perturbations or errors, allowing reliable learning despite noise at every logical layer of a given quantum learning protocol. Clarifying the relationship between experimental learnability and the robustness of (macroscopic) observables to local errors remains an open question.
@@
Finally, [huang2022foundations] develops a theoretical framework for quantum learning with noisy quantum devices, in a setting distinct from ours. They study noisy device learning from noisy access, whereas we consider learning uncharacterized and noisy systems with a fault-tolerant quantum computer. Their work analyzes a task similar to the Dec-IP problem from Definition F.2, which we use to establish quantum advantage with noisy access. However, [huang2022foundations] considers access to perfect copies of the unknown state and diamond-error noise in the action of each gate for the upper bound; moreover, the given lower bounds neglect noise.
Quantum computation with noisy oracles.
@@
The NBQP model also characterizes quantum computational problems making calls to noisy oracles, beyond the context of learning from quantum experiments. Ref. [Cross_2015] shows that a modification of the Bernstein-Vazirani problem retains a quantum speedup when only the output of the oracle is corrupted by depolarizing noise; it is simple to see that, under interstitial noise, this speedup vanishes.
@@
Several works study Grover’s search algorithm [grover1996fastquantummechanicalalgorithm] with various models of oracle noise, including phase error [Shenvi_2003], simple oracle call failure [regev2012impossibilityquantumspeedupfaulty, ambainis_grover], and phase inversions [long_grover], all finding that the time complexity under constant-strength noise reverts to asymptotically linear in the database size. Furthermore, [nisq] demonstrates such a slowdown for NISQ algorithms. Most aligned with our work, [rosmanis2023quantumsearchnoisyoracle] shows that if Grover oracle calls are sandwiched between layers of constant-strength global depolarizing or dephasing noise with all other operations perfect (a stronger computational model than NBQP), then the speedup once again breaks down, and in [rosmanis2024addendumquantumsearchnoisy] the result is strengthened to hold when depolarizing noise is applied only to a single qubit before and after the oracle query. While Grover oracles for database search tasks may eventually be implemented fault-tolerantly, these results suggest that achieving asymptotic metrological gains from the recently proposed Grover-enhanced sensing strategy in [allen2025quantumcomputingenhancedsensing] may be difficult when ambient noise acts on the sensor qubits, effectively implementing an error channel before each oracle query.
@@ Appendix B Definitions
For the definition of the NISQ complexity class, we will need to make reference to a noise model for quantum computation. The most convenient is constant depolarizing noise per qubit, although as emphasized in [nisq] other noise models are suitable as well. To fix notation, a single-qubit depolarizing channel will be defined as follows.
@@
In the NISQ model, each use of the quantum device consists of a single noisy circuit call $\textnormal{NQC}^{O}_{\lambda}(n^{\prime},\{U\})$ followed by a complete measurement of all qubits in the computational basis; the only memory across calls is classical. In particular, the algorithm cannot perform intermediate quantum measurements, reset a subset of qubits, or introduce fresh clean ancillas in the middle of a coherent computation. Quantum error-correction and fault-tolerance schemes rely precisely on such operations: one repeatedly measures stabilizers, discards or resets noisy ancillas, and thereby pumps entropy out of the encoded state. Since none of this is available in the NISQ model, one can prove (see e.g. [nisq]) that the effective noise on the quantum state cannot be suppressed over time, and the class is inherently incapable of implementing full fault-tolerant quantum error correction.
@@ C.1 Quantum information theory toolkit
Lemma C.2 (Simulating arbitrary POVMs with rank-1 POVMs, e.g. Lemma 4.8 in [chen2021exponentialseparationslearningquantum]).
@@
Lemma C.5 ($\mathcal{M}_{k}^{cn}$ POVMs are $c$-query $k$-qubit circuits, Section 8.1 in [weiyuan_paulis]).
@@ C.2 Tree representations for learning lower bounds
A powerful tool for proving information-theoretic lower bounds on the sample complexity of quantum learning tasks is the tree representation of a quantum learning algorithm. In this work, we bring together the learning tree formalisms proposed in previous works for memoryless algorithms, bounded-memory algorithms with limited memory lifetime, and bounded-memory algorithms with unbounded coherent lifetime [weiyuan_paulis, chen2021exponentialseparationslearningquantum, nisq]. These frameworks fall under the following definition.
@@
We will leverage learning trees to prove query lower bounds for property testing problems (such as the many-vs. one distinguishing problem) using Le Cam’s two point method [Yu1997LeCam]. The potential outcomes of a quantum learning algorithm for many-vs.-one distinguishing will be stored in classical memory, and encoded in the leaf nodes of the learning tree $\mathcal{T}$. Some leaves will correspond to outputting the maximally-mixed hypothesis, while others will guess the alternative event. If the two distributions over leaves $\ell$, namely $p_{\rho_{x}}(\ell)$ (given $\rho_{x}$ sampled from $\mathcal{D}$) and $p_{\mathds{1}/2^{n}}(\ell)$ (given the maximally mixed state) are very close to one another, our learning algorithm cannot successfully distinguish the two hypotheses with high accuracy. Formally:
@@
For a review of information-theoretic lower bound techniques including Le Cam’s method, see e.g. [Yu1997LeCam] and its application to learning trees in [chen2021exponentialseparationslearningquantum]. Our goal will be to argue that for the total variation bound in Lemma C.11 to become appreciably large, the tree depth $T$ must grow exponentially in $n$. To do so, it suffices to bound the one-sided likelihood ratio. Let us first define this ratio, and then subsequently explain why bounding it is useful for us.
@@
    (Lemma 5.4 in [chen2021exponentialseparationslearningquantum]) If $L(\ell)\geq 1-\delta$ for all $\ell\in\textnormal{leaf}(\mathcal{T})$, then we have the bound $d_{\text{\rm TV}}(\mathbb{E}_{\mathcal{D}}[p_{\rho_{x}}(\ell)],p_{\mathds{1}/2^{n}}(\ell))\leq\delta$.
@@
    (Lemma 7 in [weiyuan_paulis]) If the condition
@@
    Lemma 6 in [weiyuan_paulis] For any $\delta\in(0,1)$,
@@
The second lemma tells us that if the likelihood ratio induced by intermediate nodes of the tree sharply concentrates (indicating that a single experiment carries little information about the unknown state), the ratio over leaves is likely close to $1$, and the tree depth must then be large to distinguish the hypotheses. Finally, our third lemma allows us to bound the total variation distance by a probabilistic leaf likelihood ratio. That is, if we consider a fixed but arbitrary root-to-leaf path and argue that, with high probability the likelihood ratio on that path is close to $1$, we can utilize Lemma C.11. This approach is useful when the bound on intermediate nodes fails, which may happen when a small, nonzero number of intermediate nodes has large likelihood ratio fluctuations despite only comprising a vanishing fraction of the eventual outcomes. For a detailed overview of these “edge-based” and “path-based” approaches, see [chen2021exponentialseparationslearningquantum].